‘Bhartdarshan’ is an Internet-based travel agency wherein customers can see videos of the

cities they plan to visit. The number of hits daily is a normally distributed random variable

with a mean of 10,000 and a standard deviation of 2,400.

a. What is the probability of getting more than 12,000 hits?

b. What is the probability of getting fewer than 9,000 hits?


1
Expert's answer
2022-05-30T09:39:04-0400

a.

"P(X>12000)=1-P(Z\le\dfrac{12000-10000}{2400})"

"\approx1-P(Z\le 0.8333)\approx0.2023"

b.

"P(X<9000)=P(Z<\dfrac{9000-10000}{2400})"

"\approx P(Z\le -0.4167)\approx0.3385"